OpenStudy (anonymous):

@whpalmer4 What is the solution to g over five equals three over two?

OpenStudy (whpalmer4):

\[\frac{g}{5}=\frac{3}2\]is that the question?

OpenStudy (anonymous):

yeah

OpenStudy (whpalmer4):

Do you know how to cross-multiply?

OpenStudy (anonymous):

not really

OpenStudy (whpalmer4):

If two fractions are equal, the numerator of the first fraction times the denominator of the second fraction is equal to the denominator of the first fraction times the numerator of the second fraction. That means that \[g*2 = 5*3\]\[2g = 15\]Solve that for \(g\)

OpenStudy (anonymous):

or yeah I do know how to

OpenStudy (anonymous):

so its 30?

OpenStudy (whpalmer4):

Does 2*30 = 15?

OpenStudy (whpalmer4):

Whenever you get an answer, you should try to check it by plugging it back into the original formula to see if it makes sense. \[\frac{30}{5} = \frac{3}{2}\]Does that look correct? Heck, no! 30/5 = 6, 3/2 = 1.5.